# EC-SLAM: Real-time Dense Neural RGB-D SLAM System with Effectively Constrained Global Bundle Adjustment

### v1

@@ Abstract
We introduce EC-SLAM, a real-time dense RGB-D simultaneous localization and mapping (SLAM) system utilizing Neural Radiance Fields (NeRF). Although recent NeRF-based SLAM systems have demonstrated encouraging outcomes, they have yet to completely leverage NeRF’s capability to constrain pose optimization. By employing an effectively constrained global bundle adjustment (BA) strategy, our system makes use of NeRF’s implicit loop closure correction capability. This improves the tracking accuracy by reinforcing the constraints on the keyframes that are most pertinent to the optimized current frame. In addition, by implementing a feature-based and uniform sampling strategy that minimizes the number of ineffective constraint points for pose optimization, we mitigate the effects of random sampling in NeRF. EC-SLAM utilizes sparse parametric encodings and the truncated signed distance field (TSDF) to represent the map in order to facilitate efficient fusion, resulting in reduced model parameters and accelerated convergence velocity. A comprehensive evaluation conducted on the Replica, ScanNet, and TUM datasets showcases cutting-edge performance, including enhanced reconstruction accuracy resulting from precise pose estimation, 21 Hz run time, and tracking precision improvements of up to 50%. The source code is available at https://github.com/Lightingooo/EC-SLAM.
@@ I Introduction
Learning-based VSLAM systems have increasingly become a pivotal solution within the VSLAM domain, mainly due to their capacity for dense scene representation. Notably, VSLAM systems that utilize NeRF and its variants inputting RGB-D images have excelled in map and pose optimization via differentiable volume rendering. These systems are primarily categorized into two types based on their prediction methodologies: Multilayer Perceptron (MLP) based VSLAM [5, 6] and feature-decoder based VSLAM [1, 2, 3]. While MLP based VSLAM systems employ one or more MLPs for scene representation, offering smooth depiction, they face a challenge in pose constraints due to slow convergence. This slow convergence is attributed to the necessity of training all parameters in each iteration. Additionally, as the training scenarios increase in size, this integrated neural network struggles to address the issue of catastrophic forgetting.
@@
The contributions of this paper can be summarized as follows. Our first contribution is the development of a novel NeRF-based architecture that effectively enforces constraints within the global bundle adjustment framework. This architecture also takes advantage of the implicit loop closure capabilities of NeRF, which are manifested in the selection of sliding windows, the organization of keyframes, and the optimization of pixel sampling. This strategy greatly improves the accuracy and reliability of tracking and mapping. Furthermore, we integrate sparse parametric encodings with TSDF in order to accurately represent and optimize maps. This method allows for the quick and precise reconstruction of relevant maps. At last, we have created a reliable and immediate RGB-D dense SLAM system that takes advantage of the individual advantages of NeRF and classic SLAM systems for mapping and tracking. Our system has been proven effective through extensive experiments. Our strategy can be used as a robust benchmark for future works.
@@ II Related Work
NeRF-based Dense VSLAM. Although NeRF-based SLAM has lower tracking accuracy compared to traditional VSLAM, it is capable of real-time dense mapping and leverages this capability to utilize its implicit loop closure ability. It can mainly be divided into two broad categories: one integrates NeRF [49] with components from traditional SLAM systems [50, 51, 52], while the other solely rely on NeRF [5, 1, 3, 2, 53, 54, 55]. In the former type, NeRF-SLAM [51] combined the tracking component of DROID-SLAM [36] with its NeRF-based networks, while Orbeez-SLAM [50] integrated with modules from ORB-SLAM3 [18]. GO-SLAM [52] extended the tracking component of DROID-SLAM [36] by several key features. In the latter type, iMAP [5], and NICE-SLAM [1] used a single MLP and a coarse-to-fine feature grid to represent maps, respectively, but they suffered from slow convergence and struggled with scene details. Nicer-SLAM [54] performed mapping and tracking with monocular input instead of RGB-D input. For other RGB-D systems, Co-SLAM [3] and ESLAM [2] employed InstantNGP [56] and Tri-plane [57] as their scene representations, respectively, both utilizing TSDF prediction for quicker convergence. Nevertheless, these two systems overlooked the impact of randomness on the system: Co-SLAM [3] randomly selected the number and order of keyframes for global BA, failing to form effective global constraints for mapping and pose optimization. ESLAM [2] fixed the number of keyframes for BA, which limits the ability of the neural implicit field to perform implicit loop closure correction. Both systems employed random pixel sampling, which could not guarantee the uniformity and constraint of the sampled pixels. The randomness mentioned above results in these methods exhibiting weak robustness across datasets and significant variance in re-identification effectiveness for the same scene.
@@ III Method
III-A Robust multi-threaded Pipeline
We adopt a fast and effective strategy to determine the number and order of keyframes in the sliding window. To eliminate the impact of randomness, we use only the pose to decide the sliding window. Precisely, we first calculate the optical center distance and parallax angle between all keyframes and the current frame’s camera, then filter out keyframes with large parallax angles and sort them based on distance. Regarding the selection of keyframe quantity, to ensure the effectiveness of global BA, we set a maximum number threshold based on the number of sampled pixels. AS long as the number of keyframes in the sliding window does not exceed this threshold, we strive to include keyframes that can form loop closure constraints with the current frame, ensuring the presence of historical and the most recent frames. This approach not only forms practical global BA constraints but also fully leverages the inherent implicit loop closure capabilities of NeRF [49].
III-B Map Representation and Volume Rendering
@@
In addition to the network architecture, we utilize the One-blob [59] encoding for positional representation as referenced in [3]. We ascertain a given spatial point’s One-blob encoding and hash features, which are subsequently fed into the corresponding TSDF and color decoders. Unlike [3], we do not employ the intermediate embedding generated by the TSDF decoder when decoding color. Thus, the inputs to both the TSDF and color decoders are identical. We observed that such an approach facilitates more rapid network convergence.
Volume Rendering. We employ a uniform sampling strategy based on depth and spatial distribution to acquire points from the optical center to the object surface. Here, $N_{u}$ and $N_{d}$ represent the number of points sampled uniformly in space and based on depth. For depth-based sampling, we sample $N_{d}$ points within the interval $[d-d_{t},d+d_{t}]$, where $d$ is the object surface depth and $d_{t}$ is a hyperparameter that defines the sampling density.
@@ IV Experiments
This section demonstrates that our system outperforms current NeRF based systems on three widely-used datasets while maintaining a rapid processing speed. The ablation studies further confirm the efficacy of our approach.
@@ IV-A Experimental Setup
Baselines. We benchmark our method against four existing state-of-the-art NeRF based RGB-D dense VSLAM methods: iMAP [5], NICE-SLAM [1], ESLAM [2] and Co-SLAM [3] and three classic methods: Kintinuous [31], BAD-SLAM [61], ORB-SLAM2 [17].
@@ IV-B Experimental Results
Evaluation on Replica [7]. We evaluated the reconstruction and tracking performance of our system in eight scenes of the Replica dataset [7], with results detailed in Tab. I. Each metric consists of the mean and variance obtained by randomly running the system five times on each scene of the dataset. It can be observed that other methods exhibit significant fluctuations in accuracy across different scenes, indicating their poor robustness. This is attributed to the randomness of sampling in NeRF.Both configurations of our system outperform other methods in tracking metrics with speed of 12-21 Hz, achieving up to a 50% improvement in accuracy. For the Lite configuration, we achieved a precision similar to ESLAM [2] (the previous state-of-the-art) with an FPS of 21Hz. For the Full configuration, we improved the precision by 50% relative to ESLAM [2] , at a similar FPS of approximately 12Hz.
Regarding mapping metrics, our system also reached SOTA levels, and our model has a minimal number of parameters (Tab. IV). The qualitative reconstruction results of the Replica dataset [7] are shown in Fig. 3, where our system constructs more accurate scenes than other systems. The mapping results of other systems are either overly smooth, failing to adequately display high-frequency information, or excessively sharp, unable to represent the natural continuity of the scene. The reconstructed scenes of our system exhibit a good tradeoff between sharpness and smoothness, which is particularly evident in the reconstruction of small objects within the scenes, benefiting from our accurate pose estimation and sampling strategy.
Evaluation on ScanNet [4]. Although the Replica dataset [7] provides accurate reconstructed scenes for evaluating reconstruction metrics, as a synthetic indoor dataset, it does not fully represent the real world. Therefore, we also tested our system on the ScanNet dataset [7], a larger-scale real indoor dataset, which can exhibits the robustness of our system. The tracking results of our system on ScanNet [4] are detailed in Tab. II. Each metric consists of the mean and variance obtained by randomly running the system five times on each scene. Our system outperforms other systems in most scenes with a speed of 8 Hz, demonstrating its robustness to real-world environments. Our system also demonstrates higher stability in multiple runs, with much lower standard deviations.
@@
Evaluation on TUM RGB-D [8]. Consistent with previous literature [1, 3, 2], we further tested our system on three scenes of the TUM dataset. Each result is the average of five random runs of the system. As shown in Tab. III, our system achieved state-of-the-art results in most scenes, further demonstrating our system’s cross-dataset robustness. Meanwhile, our speed is the fastest in NeRF based SLAM. Although SLAM systems based on NeRF have lower tracking accuracy compared to traditional SLAM systems, our system significantly narrows this gap with accurate and dense map, providing a robust baseline for future work.
@@
TABLE II: Tracking Results on the Scannet dataset [4]
@@ IV-C Implicit loop closure capabilities of NeRF
TABLE III: Tracking Results on the Tum dataset [8]
@@ IV-D Runtime Analysis
TABLE IV: Time and Memory Analysis.
Method FPS Memory Param.
Replica iMAP [5] 0.4 0.20 M
NICE-SLAM [1] 0.9 17.40 M
Co-SLAM [3] 17.1 0.26 M
ESLAM [2] 12.1 9.29 M
Ours (Lite) 21.4 0.26 M
Ours (Full) 11.9 0.26 M
ScanNet iMAP [5] 0.1 0.2 M
NICE-SLAM [1] 0.7 10.3 M
Co-SLAM [3] 6.4 0.8 M
ESLAM [2] 4.1 10.5 M
Ours 8.6 0.8 M
@@
We evaluated the model size and FPS produced by all systems on the Replica [7], and ScanNet [4] datasets. For the Replica dataset [7], we tested the average parameter count and FPS for eight scenes, and for the ScanNet dataset [4], we assessed the parameter count and runtime for scene0000. The details are provided in Tab. IV. All results were obtained on the same computer configuration. The results indicate that our system operates faster than others, with a small model size as Co-SLAM [3], attributed to using a hash sparse grid representation of the map and our one-time sampling iteration optimization strategy.
@@ IV-E Ablations
We conducted ablation experiments on our effectively constrained global BA strategy and feature-based and uniform sampling, detailed in Tab. V. There are four comparative experiments: one without the effectively constrained global BA strategy, one without feature-based and uniform sampling, one with neither, and one with both. Without the effectively constrained global BA strategy, we randomly selected the number and order of keyframes in the sliding window and the number of sampled pixels per keyframe. Without feature-based and uniform sampling, we randomly sampled pixels. The rest of the parameters remained the same in all experiments except for the listed variable differences. The results show that both optimization strategies significantly reduce our tracking error and improve our mapping accuracy. It is noted that FUS contributes to a more significant improvement in system accuracy compared to EBA, as it markedly reduces randomness and fully leverages the implicit loop closure capabilities of NeRF.
TABLE V: Ablation study.

### v2

# EC-SLAM: Effectively Constrained Neural RGB-D SLAM with Sparse TSDF Encoding and Global Bundle Adjustment

@@ Abstract
We introduce EC-SLAM, a real-time dense RGB-D simultaneous localization and mapping (SLAM) system leveraging Neural Radiance Fields (NeRF). While recent NeRF-based SLAM systems have shown promising results, they have yet to fully exploit NeRF’s potential to constrain pose optimization. EC-SLAM addresses this by using sparse parametric encodings and Truncated Signed Distance Fields (TSDF) to represent the map, enabling efficient fusion, reducing model parameters, and accelerating convergence. Our system also employs a globally constrained Bundle Adjustment (BA) strategy that capitalizes on NeRF’s implicit loop closure correction capability, improving tracking accuracy by reinforcing constraints on keyframes most relevant to the current optimized frame. Furthermore, by integrating a feature-based and uniform sampling strategy that minimizes ineffective constraint points for pose optimization, we reduce the impact of random sampling in NeRF. Extensive evaluations on the Replica, ScanNet, and TUM datasets demonstrate state-of-the-art performance, with precise tracking and reconstruction accuracy achieved alongside real-time operation at up to 21 Hz. The source code is available at https://github.com/Lightingooo/EC-SLAM.
@@ I Introduction
Learning-based VSLAM systems have increasingly become a pivotal solution within the VSLAM domain, mainly due to their capacity for dense scene representation. Notably, VSLAM systems that utilize NeRF [5] and its variants inputting RGB-D images have excelled in map and pose optimization via differentiable volume rendering. These systems are primarily categorized into two types based on their prediction methodologies: Multilayer Perceptron (MLP) based VSLAM [6, 7] and feature-decoder based VSLAM [1, 2, 3]. While MLP based VSLAM systems employ one or more MLPs for scene representation, offering smooth depiction, they face a challenge in pose constraints due to slow convergence. This slow convergence is attributed to the necessity of training all parameters in each iteration. Additionally, as the training scenarios increase in size, this integrated neural network struggles to address the issue of catastrophic forgetting.
@@
The contributions of this paper can be summarized as follows. Firstly, we integrate sparse parametric encodings with TSDF in order to accurately represent and optimize maps. This method allows for the quick and precise reconstruction of relevant maps. Our second contribution is the development of a novel NeRF-based architecture that effectively enforces constraints within the global bundle adjustment framework. This architecture also takes advantage of the implicit loop closure capabilities of NeRF, which are manifested in the selection of sliding windows, the organization of keyframes, and the optimization of pixel sampling. This strategy greatly improves the accuracy and reliability of tracking and mapping. At last, we have created a reliable and immediate RGB-D dense SLAM system that takes advantage of the individual advantages of NeRF and classic SLAM systems for mapping and tracking. Our system has been proven effective through extensive experiments. Our strategy can be used as a robust benchmark for future works.
@@ II Related Work
NeRF-based Dense VSLAM. Although NeRF-based SLAM has lower tracking accuracy compared to traditional VSLAM, it is capable of real-time dense mapping and leverages this capability to utilize its implicit loop closure ability. It can mainly be divided into two broad categories: one integrates NeRF [5] with components from traditional SLAM systems [50, 51, 52], while the other solely rely on NeRF [6, 1, 3, 2, 53, 54, 55, 56, 57, 58].
In the former type, NeRF-SLAM [51] combined the tracking component of DROID-SLAM [37] with its NeRF-based networks, while Orbeez-SLAM [50] integrated with modules from ORB-SLAM3 [19]. GO-SLAM [52] extended the tracking component of DROID-SLAM [37] by several key features. In the latter type, iMAP [6], and NICE-SLAM [1] used a single MLP and a coarse-to-fine feature grid to represent maps, respectively, but they suffered from slow convergence and struggled with scene details. Nicer-SLAM [54] performed mapping and tracking with monocular input instead of RGB-D input. For other RGB-D systems, Co-SLAM [3] and ESLAM [2] employed InstantNGP [59] and Tri-plane [60] as their scene representations, respectively, both utilizing TSDF prediction for quicker convergence. In addition to the systems mentioned above, several SLAM systems [7, 61, 62] have integrated semantic information, which enriches the generated maps but also increases resource consumption. Furthermore, with the rise of 3D Gaussian Splatting (3DGS)  [63], many SLAM systems [64, 65] have extended 3DGS into the SLAM domain, leading to high-quality rendering. However, current 3DGS-based systems still struggle with accurate surface reconstruction and tend to run slowly, falling short of the real-time requirements expected of SLAM systems.
@@ III Method
III-A Map Representation and Volume Rendering
@@
In addition to the network architecture, we utilize the One-blob encoding [3, 66] for positional representation. We ascertain a given spatial point’s One-blob encoding and hash features, which are subsequently fed into the corresponding TSDF and color decoders. Unlike [3], we do not employ the intermediate embedding generated by the TSDF decoder when decoding color. Thus, the inputs to both the TSDF and color decoders are identical. We observed that such an approach facilitates more rapid network convergence.
Volume Rendering. We employ a uniform sampling strategy based on depth and spatial distribution to acquire points from the optical center to the object surface. In conjunction with the TSDF representation, we perform uniform sampling along the spatial ray and on the surface of the object, resulting in a total of $N=N_{u}+N_{d}$ sampled points in the space. Here, $N_{u}$ and $N_{d}$ represent the number of points sampled uniformly in space and based on depth. For depth-based sampling, we sample $N_{d}$ points within the interval $[d-d_{t},d+d_{t}]$, where $d$ is the object surface depth and $d_{t}$ is a hyperparameter that defines the sampling density.
@@
III-B Robust multi-threaded Pipeline
We adopt a simple, fast ,and effective strategy to determine the number and order of keyframes in the sliding window. To eliminate the impact of randomness, we use only the pose to decide the sliding window. Precisely, we first calculate the optical center distance $d$ and parallax angle $r$ between all keyframes and the current frame’s camera, then filter out keyframes with large parallax angles and sort them based on distance.
Regarding the selection of keyframe quantity, to ensure the effectiveness of global BA, we set a maximum number threshold based on the number of sampled pixels. AS long as the number of keyframes in the sliding window does not exceed this threshold, we strive to include keyframes that can form loop closure constraints with the current frame, ensuring the presence of historical and the most recent frames. This approach not only forms practical global BA constraints but also fully leverages the inherent implicit loop closure capabilities of NeRF [5].
@@ IV Experiments
This section demonstrates that our system outperforms current NeRF and 3DGS based systems on three widely-used datasets while maintaining a rapid processing speed. The ablation studies further confirm the efficacy of our approach.
@@ IV-A Experimental Setup
Baselines. We benchmark our method against seven existing state-of-the-art NeRF based RGB-D dense VSLAM methods: iMAP [6], NICE-SLAM [1], ESLAM [2], Co-SLAM [3], Loopy-SLAM [57], SNI-SLAM [61], NIS-SLAM [62], two 3DGS-based methods: Gaussian Splatting SLAM [64], SplaTAM [65] and three classic methods: Kintinuous [32], BAD-SLAM [69], ORB-SLAM2 [18].
For NeRF-based methods, we note that Loopy-SLAM [57] is neural point based method, which can get high reconstruction results but operates at a slow speed. SNI-SLAM [61], NIS-SLAM [62] are semantic NeRF based SLAM systems. Besides, 3DGS based methods are hard to extract mesh since the surface generated from these methods is ambiguous. For a fair comparison, we only compare the tracking results with these systems.
@@ IV-B Experimental Results
Evaluation on Replica [8]. We evaluated the reconstruction and tracking performance of our system in eight scenes of the Replica dataset [8], with results detailed in Tab. I and Tab. II. Each metric consists of the mean and variance obtained by randomly running the system five times on each scene of the dataset. It can be observed that other methods exhibit significant fluctuations in accuracy across different scenes, indicating their poor robustness. This is attributed to the randomness of sampling in NeRF. Both configurations of our system outperform other methods among most scenes with speed of 12-21 Hz. For the Lite configuration, we achieved a precision similar to ESLAM [2] (the previous state-of-the-art) with an FPS of 21Hz. For the Full configuration, we improved the precision by 50% relative to ESLAM [2] , at a similar FPS of approximately 12Hz.
Regarding mapping metrics, our system also reached SOTA levels, and our model has a minimal number of parameters (Tab. VI). The qualitative reconstruction results of the Replica dataset [8] are shown in Fig. 3, where our system constructs more accurate scenes than other systems. The mapping results of other systems are either overly smooth, failing to adequately display high-frequency information, or excessively sharp, unable to represent the natural continuity of the scene. The reconstructed scenes of our system exhibit a good tradeoff between sharpness and smoothness, which is particularly evident in the reconstruction of small objects within the scenes, benefiting from our accurate pose estimation and sampling strategy.
TABLE II: Tracking Results on 3DGS-based and NeRF-based Systems
Methods Replica FPS
r0 r1 r2 o0 o1 o2 o3 o4 Avg.
SNI [61] 0.33 0.41 0.32 0.62 0.47 0.50 0.55 0.45 0.46 2.2
NIS [62] 0.30 0.40 0.36 0.29 0.31 0.92 0.67 0.44 0.46 5.0
GSS [64] 0.76 0.37 0.23 0.66 0.72 0.30 0.19 1.46 0.58 0.7
SplaTAM [65] 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36 0.2
Loopy [57] 0.24 0.24 0.28 0.26 0.40 0.29 0.22 0.35 0.29 0.5
Ours(Full) 0.27 0.28 0.32 0.24 0.25 0.31 0.38 0.30 0.29 11.9
  • Compared to the tracking results of other NeRF and 3DGS systems. Different From Tab. I, we only provide the tracking results here, as some systems are either not open-sourced, not suitable for comparison, or do not provide additional metrics.
Evaluation on ScanNet [4]. Although the Replica dataset [8] provides accurate reconstructed scenes for evaluating reconstruction metrics, as a synthetic indoor dataset, it does not fully represent the real world. Therefore, we also tested our system on the ScanNet dataset [8], a larger-scale real indoor dataset, which can exhibits the robustness of our system. The tracking results of our system on ScanNet [4] are detailed in Tab. III and Tab. IV. Each metric consists of the mean and variance obtained by randomly running the system five times on each scene. Our system outperforms other systems in most scenes with a speed of 8 Hz, demonstrating its robustness to real-world environments. Our system also demonstrates higher stability in multiple runs, with much lower standard deviations.
@@
TABLE III: Tracking Results of Scannet Dataset.
Methods $0$ 0059 0106 0169 0181 0207 Avg. FPS
SNI [61] 6.9 7.4 7.2 - - 11.9 - -
NIS [62] 8.7 9.6 8.4 5.6 - 7.1 - -
GSS [64] 9.6 6.2 7.1 10.7 18.2 7.5 9.8 2.3
SplaTAM [65] 12.8 10.1 17.7 12.1 11.1 7.5 11.9 0.2
Loopy [57] 4.2 7.5 8.3 7.5 10.6 7.9 7.7 0.5
Ours 5.3 8.2 7.3 5.1 8.9 7.0 6.9 8.6
  • Compared to the tracking results of other NeRF and 3DGS systems. Different From Tab. IV, we only provide the RMSE results here, as some systems are not open-sourced. -- indicates that the system is not open-sourced and no results are available.
Evaluation on TUM RGB-D [9]. Consistent with previous literature [1, 3, 2], we further tested our system on three scenes of the TUM dataset. Each result is the average of five random runs of the system. As shown in Tab. V, our system achieved state-of-the-art results in most scenes, further demonstrating our system’s cross-dataset robustness. Meanwhile, our speed is the fastest in NeRF based SLAM. Although SLAM systems based on NeRF have lower tracking accuracy compared to traditional SLAM systems, our system significantly narrows this gap with accurate and dense map, providing a robust baseline for future work.
@@
TABLE IV: Tracking Results on the Scannet dataset [4]
@@ IV-C Implicit loop closure capabilities of NeRF
TABLE V: Tracking Results on the Tum dataset [9]
@@ IV-D Runtime Analysis
TABLE VI: Time and Memory Analysis.
Method FPS Model Param.
Replica NICE-SLAM [1] 0.9 49.0 M
Co-SLAM [3] 17.1 7.9 M
ESLAM [2] 12.1 23.2 M
GSS [64] 0.7 24.2 M
SplaTAM [65] 0.2 243.4 M
Ours (Lite) 21.4 8.2 M
Ours (Full) 11.9 8.2 M
ScanNet NICE-SLAM [1] 0.7 88.7 M
Co-SLAM [3] 6.4 9.9 M
ESLAM [2] 4.1 70.9 M
GSS [64] 2.3 5.6 M
SplaTAM [65] 0.2 156.4 M
Ours 8.6 11.3 M
@@
We evaluated the model size and FPS produced by all systems on the Replica [8], and ScanNet [4] datasets. For the Replica dataset [8], we tested the parameter count and FPS for room0, and for the ScanNet dataset [4], we assessed the parameter count and runtime for scene0000. The details are provided in Tab. VI. All results were obtained on the same computer configuration. The results indicate that our system operates faster than others, with a small model size as Co-SLAM [3], attributed to using a hash sparse grid representation of the map and our one-time sampling iteration optimization strategy.
@@ IV-E Ablations
We conducted ablation experiments on our effectively constrained global BA strategy and feature-based and uniform sampling, detailed in Tab. VII. There are four comparative experiments: one without the effectively constrained global BA strategy, one without feature-based and uniform sampling, one with neither, and one with both. Without the effectively constrained global BA strategy, we randomly selected the number and order of keyframes in the sliding window and the number of sampled pixels per keyframe. Without feature-based and uniform sampling, we randomly sampled pixels. The rest of the parameters remained the same in all experiments except for the listed variable differences. The results show that both optimization strategies significantly reduce our tracking error and improve our mapping accuracy. It is noted that FUS contributes to a more significant improvement in system accuracy compared to EBA, as it markedly reduces randomness and fully leverages the implicit loop closure capabilities of NeRF.
TABLE VII: Ablation study.